Hallc beamtest runlist highrate » History » Version 1
Jixie Zhang, 01/23/2023 11:11 AM
| 1 | 1 | Jixie Zhang | h1. Hallc beamtest runlist highrate |
|---|---|---|---|
| 2 | |||
| 3 | --------------------------------------------------- |
||
| 4 | List of HVs: |
||
| 5 | board1, 6kV, 8-ch: |
||
| 6 | Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D |
||
| 7 | HV:_______0,_____0,_____0,_____0,_0,_0,_0,__1000 |
||
| 8 | |||
| 9 | board2, 2kV, 16-ch: |
||
| 10 | Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC |
||
| 11 | HV:___1200,_1600,_1650,_0,_____850,____1000,_1600,___________1515,___________1800,___________1730,_____________930,_____________950,_____________850,__1000,__1000,__1000 |
||
| 12 | |||
| 13 | --------------------------------------------------- |
||
| 14 | |||
| 15 | 3944 By Jixie: debug FADC trigger |
||
| 16 | |||
| 17 | 3949 cosmic(6h) and tune beam no target(3h). By Jixie: debug FADC trigger |
||
| 18 | |||
| 19 | 3950 Pedestal. No HV |
||
| 20 | 3951 cosmic, no beam. debug FADC trigger |
||
| 21 | 3952 debug gem ssp |
||
| 22 | 3953 cosmic, no beam. debug FADC trigger |
||
| 23 | 3954 production. 21:16 Jan 11st - 23:14 Jan 11st |
||
| 24 | daq rate = 12.5kHz |
||
| 25 | |||
| 26 | 3955 production. 22:16 Jan 11st - 23:32 Jan 11st |
||
| 27 | daq rate = 12kHz |
||
| 28 | |||
| 29 | |||
| 30 | ==change prescale factors, disable internal pulse trigger for the first 100k events == |
||
| 31 | |||
| 32 | /*Set prescale for each TS#*/ |
||
| 33 | tiSetInputPrescale(1,0); |
||
| 34 | tiSetInputPrescale(2,0); |
||
| 35 | tiSetInputPrescale(3,5); |
||
| 36 | tiSetInputPrescale(4,3); |
||
| 37 | tiSetInputPrescale(5,3); |
||
| 38 | tiSetInputPrescale(6,0); |
||
| 39 | |||
| 40 | ==use hallc_fadc_ssp config== |
||
| 41 | |||
| 42 | 3956 trash. 22:34 Jan 11st |
||
| 43 | tried to use hallc_fadc_ssp config, could not start CODA |
||
| 44 | |||
| 45 | ==use hallc_fadc config== |
||
| 46 | |||
| 47 | 3957 production. 23:37 Jan 11st - 23:45 Jan 11st |
||
| 48 | daq rate = 12.5kHz @ 2uA , this is the maximum of the daq rate for uvasoild2 |
||
| 49 | Threshold: Cer_Sum(30mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV) |
||
| 50 | |||
| 51 | |||
| 52 | 3958 production. 23:48 Jan 11st - 00:47 Jan 12th |
||
| 53 | Threshold: Cer_Sum(100mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV) |
||
| 54 | |||
| 55 | ==Turn off HVs== |
||
| 56 | |||
| 57 | Jan 12th, 12:30 |
||
| 58 | |||
| 59 | 3959-3971 Trash, deug daq |
||
| 60 | |||
| 61 | |||
| 62 | ----------------------------------------------------------------------- |
||
| 63 | ==Turn on HVs, all Cer HVs are changed from 1000V to 950V, Shower_T(SDU#5) lower from 850V to 800V== |
||
| 64 | |||
| 65 | --------------------------------------------------- |
||
| 66 | List of HVs: |
||
| 67 | board1, 6kV, 8-ch: |
||
| 68 | Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D |
||
| 69 | HV:_______0,_____0,_____0,_____0,_0,_0,_0,__950 |
||
| 70 | |||
| 71 | board2, 2kV, 16-ch: |
||
| 72 | Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC |
||
| 73 | HV:___1200,_1600,_1650,_0,_____850,____1000,_1600,___________1515,___________1800,___________1730,_____________930,_____________950,_____________800,___950,___950,___950 |
||
| 74 | |||
| 75 | --------------------------------------------------- |
||
| 76 | |||
| 77 | only TS#6 is enabled, PS#(15,15,15,15,15,15,0) |
||
| 78 | Threshold: Cer_Sum(300mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV) |
||
| 79 | |||
| 80 | |||
| 81 | Cer_HV = 950V |
||
| 82 | 3972 pulse, no beam. 18:11 Jan 12th - 18:18 Jan 12th, Cer_HV = 950V |
||
| 83 | 3973 pulse, no beam. 18:19 Jan 12th - 18:24 Jan 12th, Cer_HV = 950V |
||
| 84 | 15.5kHz @ no beam, pulse trigger |
||
| 85 | |||
| 86 | Cer_HV = 970V |
||
| 87 | 3974 pulse, no beam. 18:26 Jan 12th - 18:32 Jan 12th, Cer_HV = 970V |
||
| 88 | 3975 pulse, no beam. 18:32 Jan 12th - 18:52 Jan 12th, Cer_HV = 970V |
||
| 89 | |||
| 90 | Cer_HV = 930V |
||
| 91 | 3976 pulse, no beam. 18:54 Jan 12th - 19:01 Jan 12th, Cer_HV = 930V |
||
| 92 | 3977 pulse, no beam. 19:03 Jan 12th - 19:09 Jan 12th, Cer_HV = 930V |
||
| 93 | |||
| 94 | Cer_HV = 1000V |
||
| 95 | 3978 pulse, no beam. 19:54 Jan 12th - 20:06 Jan 12th, Cer_HV = 1000V |
||
| 96 | 3979 pulse, no beam. 20:06 Jan 12th - 20:16 Jan 12th, Cer_HV = 1000V |
||
| 97 | |||
| 98 | Cer_HV = 1050V |
||
| 99 | 3980 pulse, no beam. 20:17 Jan 12th - 20:24 Jan 12th, Cer_HV = 1050V |
||
| 100 | 3981 pulse, no beam. 20:25 Jan 12th - 20:39 Jan 12th, Cer_HV = 1050V |
||
| 101 | ----------------------------------------------------------------------- |
||
| 102 | |||
| 103 | 3982 cosmic production, |
||
| 104 | first 100k is pulse trigger, 22:36 - 23:28 Jan 12th |
||
| 105 | debug TS#6, all trigger bits enabled, PS#(0,0,0,0,0,0,0) |
||
| 106 | Threshold: Cer_Sum(300mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV) |
||
| 107 | |||
| 108 | |||
| 109 | ==cosmic test for Shower only, HVs for others are off, first 100k is pulse trigger== |
||
| 110 | TS#4 only PS#(15,15,15,0,15,15) |
||
| 111 | Threshold: Cer_Sum(300mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV) |
||
| 112 | |||
| 113 | 3983 cosmic production, 23:34 - 00:25 Jan 13th |
||
| 114 | Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5) |
||
| 115 | HV:_______________930,_____________950,_____________800 |
||
| 116 | |||
| 117 | ==00:27 Jan 13th, turn all HVs off== |
||
| 118 | |||
| 119 | 3984 cosmic production, 08:25 Jan 13th - 13:03 Jan 13th |
||
| 120 | Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5) |
||
| 121 | HV:______________1030,____________1050,_____________850 |
||
| 122 | |||
| 123 | 3985 cosmic production, 13:04 Jan 13th - 17:13 Jan 13th |
||
| 124 | Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5) |
||
| 125 | HV:______________1030,____________1050,_____________850 |
||
| 126 | see 12 minutes of beam ... detectors are fully saturated |
||
| 127 | |||
| 128 | 3986 cosmic production, 17:13 Jan 13th - 18:32 Jan 13th |
||
| 129 | Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5) |
||
| 130 | HV:______________1030,____________1050,_____________850 |
||
| 131 | see 5 minutes of beam ... detectors are fully saturated |
||
| 132 | |||
| 133 | ==all HVs are off== |
||
| 134 | all 6 trigger bits are on, no prescale factors |
||
| 135 | |||
| 136 | ==dark current run (beam on but HV off, pulse trigger)== |
||
| 137 | 3987 cosmic production, 18:36 Jan 13th - 18:42 Jan 13th |
||
| 138 | I tried to get beam on data, but beam is off |
||
| 139 | |||
| 140 | |||
| 141 | 3988 dark current run, 11:05 Jan 14th - 11:11 Jan 14th |
||
| 142 | beam current 40uA, HV off, pulse trigger |
||
| 143 | |||
| 144 | 3989 dark current run, 11:12 Jan 14th - 11:20 Jan 14th |
||
| 145 | beam current 40uA, HV off, pulse trigger |
||
| 146 | |||
| 147 | ======================================================================= |
||
| 148 | ==5uA to test the prescale factors== |
||
| 149 | HV for CerD is off (forgot to turn on) |
||
| 150 | |||
| 151 | Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC |
||
| 152 | HV:___1200,_1600,_1650,_0,_____850,____1000,_1600,___________1515,___________1800,___________1730,____________1030,_____________1050,____________850,___950,___950,___950 |
||
| 153 | |||
| 154 | |||
| 155 | 3990 all PS=0 TS1-6 daq rate=15.5kHz |
||
| 156 | 3991 all PS=0 TS1-6 daq rate=15.5kHz |
||
| 157 | 3992 PS4=5, TS1-6 daq rate=15.5kHz |
||
| 158 | 3993 TS4 only PS4=5 daq rate=12.5kHz |
||
| 159 | |||
| 160 | ======================================================================= |
||
| 161 | Increase shower_T HV from 850 to 900 |
||
| 162 | Decrease shower_R HV from 1050 to 1030 |
||
| 163 | HVs for SC_D, SC_A, LASPD_T,LASPD_B, PreSh_# were all got lower |
||
| 164 | |||
| 165 | Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC |
||
| 166 | HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___930,___930,___930 |
||
| 167 | |||
| 168 | CerD HV is also 930V |
||
| 169 | |||
| 170 | 3994-3995 trash, HV boardB tripped |
||
| 171 | |||
| 172 | 3996 dark current run, beam off, HVs all on, pulse trigger, 930V |
||
| 173 | 3997 dark current run, beam off, HVs all on, pulse trigger, 930V |
||
| 174 | |||
| 175 | 3998 dark current run, beam off, HVs all on, pulse trigger, 940V |
||
| 176 | 3999 dark current run, beam off, HVs all on, pulse trigger, 940V |
||
| 177 | |||
| 178 | 4000 dark current run, beam off, HVs all on, pulse trigger, 920V |
||
| 179 | 4001 dark current run, beam off, HVs all on, pulse trigger, 920V |
||
| 180 | |||
| 181 | 4002 dark current run, beam off, HVs all on, pulse trigger, 910V |
||
| 182 | 4003 dark current run, beam off, HVs all on, pulse trigger, 910V |
||
| 183 | |||
| 184 | 4004 dark current run, beam off, HVs all on, pulse trigger, 950V |
||
| 185 | 4005 dark current run, beam off, HVs all on, pulse trigger, 950V |
||
| 186 | |||
| 187 | 4006 dark current run, beam off, HVs all on, pulse trigger, 900V |
||
| 188 | 4007 dark current run, beam off, HVs all on, pulse trigger, 900V |
||
| 189 | |||
| 190 | 4008 dark current run, beam off, HVs all on, pulse trigger, 960V |
||
| 191 | 4009 dark current run, beam off, HVs all on, pulse trigger, 960V |
||
| 192 | |||
| 193 | |||
| 194 | ==================================================================== |
||
| 195 | Cer HVs go back to 950V, 1uA CW beam on He3 target |
||
| 196 | |||
| 197 | 4010 beam on, but only pulse trigger |
||
| 198 | 4011 first 100k is pulse trigger, 50Hz rate without beam. TS1 to 6 are all on, PS4=5; NO real beam on this run |
||
| 199 | 4012 disabled pulse trigger. All TS#. PS4=5, no beam at |
||
| 200 | 4013 disabled pulse trigger. All TS#. PS4=9, daq rate = 13.5kHz,1uA beam |
||
| 201 | 4014 disabled pulse trigger. All TS#. PS4=11, daq rate = 13.5kHz,1uA beam |
||
| 202 | 4015 disabled pulse trigger. All TS#. PS4=15, PS5=15, daq rate = 13.5kHz,1uA beam |
||
| 203 | 4016 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS2=PS3=PS6=0, daq rate = 13.5kHz, 1uA beam |
||
| 204 | 4017 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=5, no beam |
||
| 205 | 4018 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=5, 2uA beam. 12.8KHz |
||
| 206 | 4019 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=10, 2uA beam. 8.5KHz |
||
| 207 | 4020 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=12, 2uA beam. 2.7KHz |
||
| 208 | 4021 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS3=PS6=15, PS1=0,PS2=12, 2uA beam. 11KHz |
||
| 209 | 4022 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS3=PS6=15, PS1=5,PS2=12, 1.8uA beam. 5.7KHz |
||
| 210 | |||
| 211 | Conclusion: on He3 target, PS1=5 will have 1.8kHz/uA for TS#1, PS2=12 will have 1.4kHz/uA for TS#2. |
||
| 212 | We still need to determine PS for PS4 and PS5. |
||
| 213 | After analyze run 4022, I found that for PS3=15, TS#3 has 28Hz/2uA. For PS6=15, TS#6 has 8Hz/2uA. Scaling PS from |
||
| 214 | 15 to 10, we got TS#3 rate as 450Hz/uA and TS#6 rate 130Hz/uA. |
||
| 215 | |||
| 216 | |||
| 217 | =================================================== |
||
| 218 | pedestal runs, beam off, pulse trigger, HV scan |
||
| 219 | |||
| 220 | Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC |
||
| 221 | HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___930,___930,___930 |
||
| 222 | |||
| 223 | For Cherenkov: ch13,ch14,ch15, E08C2_ch07: 910, 930, 950, 970 |
||
| 224 | For Shower: ch10,ch11: 900, 950, 1000, 1050 |
||
| 225 | For Shower: ch12: 800, 850, 900, 950 |
||
| 226 | |||
| 227 | ch 10 11 12 13 14 15 07 |
||
| 228 | 4023, 4024: HV(ch10-15,E08C2_ch07) : 800, 900, 900, 910, 910, 910, 910 |
||
| 229 | 4025, 4026: 950, 950, 850, 930, 930, 930, 930 |
||
| 230 | 4027, 4028: 1000,1000, 900, 950, 950, 950, 950 |
||
| 231 | 4029, 4030: 1050,1050, 950, 970, 970, 970, 970 |
||
| 232 | 4031, 4032: 1030,1030, 900, 950, 950, 950, 950 |
||
| 233 | |||
| 234 | =================================================== |
||
| 235 | |||
| 236 | Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC |
||
| 237 | HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___930,___930,___930 |
||
| 238 | |||
| 239 | ==turn on FADC threshold to remove no pulse events == |
||
| 240 | slot 7: FADC250_ALLCH_TET 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31 |
||
| 241 | slot 8: FADC250_ALLCH_TET 10 10 10 20 10 10 10 5 5 5 5 5 5 15 15 25 |
||
| 242 | |||
| 243 | 4033 debug FADC config, no beam, HV on. 5M pulse trigger |
||
| 244 | 4034 debug FADC config, no beam, HV on. 5M pulse trigger |
||
| 245 | |||
| 246 | ==turn on slot 8 FADC threshold to remove no pulse events == |
||
| 247 | slot 7: FADC250_ALLCH_TET 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31 |
||
| 248 | slot 8: FADC250_ALLCH_TET 10 10 10 20 10 10 10 10 10 10 10 10 10 20 20 25 |
||
| 249 | 4035 trash |
||
| 250 | 4036 debug FADC config, no beam, HV on. 5M pulse trigger |
||
| 251 | |||
| 252 | ============== |
||
| 253 | 4037 disabled pulse trigger, cosmic only, beam off HV on |
||
| 254 | |||
| 255 | ============== |
||
| 256 | 4038-4040 trash, check daq, adkust threshold |
||
| 257 | |||
| 258 | ======================================================== |
||
| 259 | ==20230123 threshold== |
||
| 260 | Threshold is applied to the height of the waveform: |
||
| 261 | TS1: CerSum(200mv or 410 adc unit) & ShowerSum(70mv or 143 adc unit); |
||
| 262 | TS2: SC_D(20mv or 164 adc unit) & ShowerSum(70mv or 143 adc unit) & SC_B(15mv or 123 adc unit). |
||
| 263 | TS3: 2 out of these 3 inputs: CerSum(200mv or 410 adc unit), ShowerSum(50mv or 103 adc unit), SC_D(20mv or 164 adc unit); |
||
| 264 | TS4: ShowerSum(100mv or 205 adc unit) |
||
| 265 | TS5: SC_B(20mv or 164 adc unit). |
||
| 266 | TS6: CerA#,CerB#,CerC#,CerD#(15mv or 31 adc unit) for each Cherenkov individual input |
||
| 267 | |||
| 268 | 5uA tune beam on LHe4, |
||
| 269 | 4041 5uA tune beam on LHe4, PS=10,10,10,10,10,10, rate=250Hz |
||
| 270 | 4042 5uA tune beam on LHe4, PS=4,4,3,1,0,0, rate=2.4kHz |
||
| 271 | 4043 5uA tune beam on LHe4, PS=3,3,2,0,0,0, rate=4kHz |
||
| 272 | |||
| 273 | 2uA CW beam on LHe4, |
||
| 274 | 4044 all TS1-6; PS=7,7,5,5,3,0, 11.5kHz, daq saturated |
||
| 275 | 4045 all TS1-6; PS=8,8,5,3,4,0, 11.2kHz, daq saturated |
||
| 276 | 4046 all TS1-6; PS=8,8,5,3,3,1, 11.2kHz, daq saturated |
||
| 277 | 4045 all TS1-6; PS=9,9,5,3,3,0, 11.2kHz, daq saturated |
||
| 278 | |||
| 279 | 2uA CW beanm on LHe4, |
||
| 280 | 4048 TS#4 only, PS4=3, rate=10.0kHz |
||
| 281 | 4049-4050 trash |
||
| 282 | 4051 TS#4 only, PS4=4, rate=5.9kHz |
||
| 283 | 4052-4054 trash, TS#5 only, no rate, TS#5 cable connection might have problem |
||
| 284 | 4055 TS#4 only, PS4=4, rate=5.6kHz |
||
| 285 | ==turn off sw test pulse in diaggui, this test inserts 100Hz pulse into FADC== |
||
| 286 | ==do not know how it affects the above data== |
||
| 287 | 4056 trash, TS#6 only, no rate, |
||
| 288 | 4057 TS#3 only, PS3=5, rate=2.0kHz |
||
| 289 | 4058 TS#3 only, PS3=3, rate=6.8kHz |
||
| 290 | 4059 trash |
||
| 291 | 4060 TS#2 only, PS2=10, rate=1.1kHz |
||
| 292 | 4061 TS#2 only, PS2= 7, rate=7.6kHz |
||
| 293 | 4062 TS#1 only, PS2= 9, rate=2.8kHz |
||
| 294 | 4063 TS#1 only, PS2= 7, rate=9.2kHz |
||
| 295 | 4064 All TS, PS=9,9,5,6,6,1, rate=7.8kHz, |
||
| 296 | |||
| 297 | 4065 All TS, PS=9,9,5,6,6,1, 20uA beam on LHe4, rate=11.2kHz, daq saturated |
||
| 298 | |||
| 299 | ======================================================== |